# Sustainable Transparency on Recommender Systems: Bayesian Ranking of Images for Explainability

### v2

@@ Abstract
[…]
@@ 1 Introduction
In recent years, the field of Artificial Intelligence (AI) has witnessed a significant rise in the development, research, and deployment of complex AI systems that can have a direct impact on users’ and citizens’ lives. From decision-making processes to personalized recommendations, they have become necessary components of numerous applications and fields, enhancing user experience and improving the communicative capabilities of institutions and companies. However, as these systems grow in complexity and influence on society, there is a pressing need to ensure their transparency and explainability, particularly in domains where user trust is paramount doshi2017accountability .
@@
The growing importance of explainability in AI systems during the past decade stems from the need to address the “black box” nature of many complex models ribeiro2016should , as users often find it challenging to comprehend why an AI system produces a specific output, such as an item or product recommendation provided by a Recommender System. This lack of transparency has proved to lead to a diminished user experience, reduced trust from society in AI systems, and, in the worst case, potentially harmful outcomes hamm2023explanation . This has popularly led to regulatory efforts to ensure transparency and accountability in AI systems. For instance, the European Union’s General Data Protection Regulation (GDPR) mandates the right to explanation in data-treating systems, granting users insight into any automated decisions made by them gdpr . Ethical-oriented guidelines, such as those published by the EU’s High-Level Expert Group on Artificial Intelligence, emphasize transparency and accountability in AI eu-guidelines . Additionally, the U.S. Federal Trade Commission (FTC) emphasized the need for explainable AI to prevent deceptive practices in a recent report ftc-report .
@@
With respect to recommendations, the increasing reliance of businesses and users on AI-based Recommender Systems (RS) means that understanding and interpreting the recommendations provided to users has become crucial. Users expect personalized recommendations that align with their preferences, but they also desire explanations for why certain recommendations are made, e.g. due to a lack of quality of the recommendation or mistrust in the system; as such, explaining the outputs of RS not only enhances user trust, but also further enables users to make informed decisions, promoting a more interactive and engaging user experience zhang2020explainable .
@@
In the quest for explainability, one promising avenue is leveraging the visual content associated with AI inputs and outputs. Visual information, such as images, plays a vital role in human perception and decision-making processes, having been explored before as a tool for explainability in other tasks zeiler2014visualizing . In the context of user-to-item recommendations, among other examples, Netflix uses a small selection of movie thumbnails crafted by experts to cater to different types of users netflix-research-2023 , and Chen, Xu, et al. chen2019personalized leverage user tastes from textual fashion reviews to highlight different image sections of the fashion item being recommended. Consequently, utilizing visual content as a means to explain recommendations can greatly enhance user understanding and satisfaction hendricks2016generating .
@@
Within the idea of visual explanations, our particular approach of interest is the use of user-generated images to provide the explanations, avoiding a need for auxiliary information (e.g. text) to distil user tastes or a manual selection of creation of the images, thus yielding more organic explanations and improving the self-sustainability and scalability of the explainable methodology.
@@
Leveraging user-created images is a promising but particularly challenging task: in the case of textual contents, a textual explanation may be associated with multiple users or items (e.g. the phrase “great visual effects” in movie reviews), which can help in tackling data sparsity li2023relationship . User-created visual explanations (e.g. an image taken in a restaurant), however, are unique to each user and item, making it difficult for a model to distil user tastes. This challenge has been tackled by both ELVis diez2020towards and MF-ELVis paz2022sustainable , which model and solve the explanation problem as a Learning to Rank (LTR) authorship task, predicting the best explanations of a (user, item) recommendation as a ranking of the images of the item most likely to have been authored by the user.
@@
While explainability is crucial, it should not come at the expense of sustainability and model compactness. AI and Deep Learning models are known for their high computational cost, carbon emissions and model sizes; these concerns highlight the need to balance model efficiency and explainability strubell2019energy cheng2018model . ELVis first provided explainability to an RS by leveraging user-created images, but used a Deep Learning-oriented architecture with a cost comparable to the RS itself. Although MF-Elvis optimized the architecture of the model with a simpler yet adequate Matrix Factorization approach, both models still need a costly training process and a large model size to obtain good results. We believe one potential culprit of this computational cost is the inadequacy of the training objective used by both ELVis and MF-ELVis, which model the task of ranking the best image explanations as a surrogate binary classification of image authorship, which may be far from optimal.
@@
    We introduce, design and implement BRIE, a novel model that leverages existing user-created images to provide visually interpretable explanations for Recommender System outputs. BRIE improves existing approaches by adopting a more adequate learning objective closer to the final use of the model, while maintaining a highly optimized and simple architecture.
@@ 2.1 Problem Formulation
Therefore, leveraging user-generated item images as visual explanations poses two main challenges: 1) extracting meaningful latent information from photographs, given their limited association with multiple users and items, and 2) addressing data sparsity in real-world datasets where the available set of photographs $\mathcal{P}$ may be insufficient compared to the set of users $\mathcal{U}$.
@@
As an explainability task, this problem assumes that the item $i$ being recommended is already given in a previous step by a Recommender System or any other information filtering system that matches users with items, and therefore focuses on finding a ranking function $f$ that can capture nuanced user preferences and photograph characteristics to provide adequate explanations of any given user-item pair $(u,i)$.
@@ 2.2 ELVis and MF-ELVis
ELVis and MF-ELVis are the current state-of-the-art models for visual explainability through user-created images in Recommender Systems. In this subsection, we describe the functioning of these models and provide insights into their design choices. Fig. 1 depicts in detail the network topologies of both models.
@@
Figure 1: Network topologies of ELVis and MF-ELVis. Both models take as inputs the IDs of user $u$ and photograph $p$, and output the predicted authorship $\mathbf{\hat{R}}_{up}$, i.e. how good is $p$ as an explanation of a recommendation of its associated item for user $u$. During training, both models minimize BCE loss $\mathcal{L}_{\text{BCE}}$ based on the predicted and real authorship $\mathbf{\hat{R}}_{up}$ and $\mathbf{R}_{up}$.
@@ 2.2.1 ELVis
Díez et al. diez2020towards propose in ELVis to capture the underlying patterns and preferences in the interactions between users and photographs by representing the interaction matrix $\mathbf{R}$ as two latent matrices: the matrix of latent user preferences $\mathbf{U}$ and the matrix of latent photograph features $\mathbf{V^{\prime}}$.
ELVis generates a latent user vector $\mathbf{U}_{u}\in\mathbb{R}^{d}$ based on the user’s ID, and obtains a latent photograph vector $\mathbf{V}_{p}^{\prime}\in\mathbb{R}^{d}$ by projecting the ResNet-V2 szegedy2017inception embedding $\mathbf{V}_{p}\in\mathbb{R}^{1536}$ of the original photograph into the latent feature subspace. The authors then propose to concatenate both vectors as $\mathbf{x}_{up}$, which is then passed through a multi-layer perceptron (MLP) $g$ to apply a series of non-linear transformations, such that
@@
The rationale behind ELVis’s design is based on using a learned similarity function, following the popular Neural Collaborative Filtering architecture he2017neural . The principle is that MLPs are general function approximators, so they should be -in theory- equally or more expressive than a fixed similarity function, such as a cosine similarity or dot product, and should capture more nuanced latent user preferences. However, the usage of an MLP comes at the expense of a very high computational cost overhead in both training and inference, motivating later approaches to abandon the usage of deep architectures.
@@ 2.2.2 MF-ELVis
A later publication suggests MF-ELVis paz2022sustainable to simplify the model architecture while maintaining performance. Similar to ELVis, MF-ELVis learns the latent user preference matrix $\mathbf{U}$ and item feature matrix $\mathbf{V^{\prime}}$. However, instead of an MLP, MF-ELVis uses a fixed similarity function, the inner product, to compute the similarity between the corresponding user vector $\mathbf{U}_{u}$ and photograph vector $\mathbf{V}_{p}^{\prime}$. For a given (user, photograph) pair, the preference score is calculated as
@@
The rationale for adopting the inner product in MF-ELVis is based on findings presented by Rendle et al. rendle2020neural . The study suggests that, in similarity-related tasks, the classic inner product can achieve performance comparable to an MLP while simplifying the model architecture. Leveraging the inner product operation, MF-ELVis enhanced the time and carbon emissions required to train the model by simplifying the model architecture. However, it utilized a much higher latent dimension $d$ to maintain competitiveness with ELVis, increasing the overall model size; this suggested that under the training protocols ELVis and MF-ELVis use, the model architecture could not be further simplified without losing performance, but other parts of the training pipeline, such as the modelled learning objective and the use of training data, could be suboptimal and an aspect to be improved.
@@ 2.3 Limitations of ELVis and MF-ELVis
Both ELVis and MF-ELVis address this issue by employing a uniform negative sampling and a positive sampled replication, as shown in the left half of Figure 2, yielding a static training dataset 40 times larger than the original with a balanced number of positive and negative samples. Specifically, for each positive sample $(u,i,p)\in\mathcal{D}$, 39 additional samples are generated and used during all the training: 19 extra replications of $(u,i,p)$, 10 negative triads $(u,i,p^{\prime})$ (where $p^{\prime}$ are from the same restaurant as $p$ but uploaded by a different user $u^{\prime}$), and 10 negative triads $(u,i^{\prime},p^{\prime})$ (where $p^{\prime}$ were uploaded by a different user $u^{\prime}$ in a restaurant other than $i$). In the field of Positive-Unlabeled Learning, this random selection of negative samples is considered a naive yet straightforward method bekker2020learning .
@@ 3 BRIE
In this section, we present BRIE (Bayesian Ranking of Images for Explainability), a novel model that builds upon MF-ELVis and addresses its limitations to surpass state-of-the-art performance in serving user-created visual content to explain recommendations, while drastically improving training and inference efficiency.
@@ 3.1 Modelling Explanation Ranking with BPR
State-of-the-art approaches, like ELVis and MF-ELVis, utilize binary classification as a surrogate task to model the Learning to Rank scenario required to select the most adequate visual explanation of an item for a given user. Instead, BRIE uses the Bayesian Pairwise Ranking (BPR) rendle2012bpr algorithm to learn the ranking function $f$. BPR is tailored to optimize model weights by maximizing the probability of correctly ranking positive “good” instances higher than negative “bad” instances.
The selection policy of these negative instances in BRIE differs from the one ELVis and MF-ELVis utilize for their binary classification learning, described in Section 2.3. At the start of each training epoch, BRIE constructs an “extended” training set $\mathcal{D^{+}}={(u,i,p,p_{neg})}$ by uniformly sampling, for each $(u,i,p)\in\mathcal{D}$, a photograph $p_{neg}$ that is not part of $\mathcal{P}_{u}\subset\mathcal{P}$, i.e. a photograph from a purely random item and not taken by user $u$. Figure 2 compares this selection policy with the one used in current state-of-the-art models. Our method is simpler than the pre-training selection used by ELVis and MF-ELVis, allows the regeneration of negative samples between epochs, and avoids storing all the pre-selected negative samples and replicated positive samples.
@@ 4.1 Datasets
The datasets used in our experiments are derived from the original publication of ELVis by Diez et al. diez2020towards , comprising photographs uploaded by users in the context of restaurant reviews on TripAdvisor, a popular hospitality establishment reviewing website. We utilize public, real-world datasets from six different cities: Gijón, Barcelona, Madrid, New York, Paris, and London, in increasing number of photographs in the dataset, providing diverse contexts for evaluation both from a socio-cultural and a data perspective.
Each sample in the raw dataset consists of a triplet $(u,i,p)\in\mathcal{D}$, where $u$ represents the user ID, $i$ represents the restaurant ID, and $p$ represents the photograph ID. For the sake of result reproducibility and comparability, we use the original train/validation/test partitions provided by the original authors.
For each test case, we have one positive sample, a photograph ($p$), that belongs to the user ($u$). The remaining photographs $p^{\prime}\in(\mathcal{P}_{i}\setminus\mathcal{P}_{u}$) of the same restaurant, contributed by different users, serve as negative samples. The goal of the model is to rank the positive sample (the user’s own photograph) as high as possible among the negative samples, indicating its relevance as an explanation. All positive samples in the test set correspond to users already present in the training set, ensuring a fair evaluation of the model’s ability to generalize to unseen user-photograph combinations.
To provide a comprehensive overview of the datasets’ characteristics, Table 1 summarizes key statistics, including the number of users, photographs, restaurants, and the overall sparsity of the datasets. We refer readers to the original ELVis publication diez2020towards for more detailed information on the dataset extraction, characteristics, and partitioning.

### v3

@@ Abstract
Recommender Systems have become crucial in the modern world, commonly guiding users towards relevant content or products, and having a large influence over the decisions of users and citizens. However, ensuring transparency and user trust in these systems remains a challenge; personalized explanations have emerged as a solution, offering justifications for recommendations. Among the existing approaches for generating personalized explanations, using existing visual content created by users is a promising option to maximize transparency and user trust. State-of-the-art models that follow this approach, despite leveraging highly optimized architectures, employ surrogate learning tasks that do not efficiently model the objective of ranking images as explanations for a given recommendation; this leads to a suboptimal training process with high computational costs that may not be reduced without affecting model performance. This work presents BRIE, a novel model where we leverage Bayesian Pairwise Ranking to enhance the training process, allowing us to consistently outperform state-of-the-art models in six real-world datasets while reducing its model size by up to 64 times and its CO2 emissions by up to 75% in training and inference.
@@ 1 Introduction
In recent years, the field of Artificial Intelligence (AI) has witnessed a significant rise in the development, research, and deployment of complex AI systems that can have a direct impact on users’ and citizens’ lives. From decision-making processes to personalized recommendations, they have become necessary components of numerous applications and fields, enhancing user experience and improving the communicative capabilities of institutions and companies. However, as these systems grow in complexity and influence on society, there is a pressing need to ensure their transparency and explainability, particularly in domains where user trust is paramount [1].
@@
The growing importance of explainability in AI systems during the past decade stems from the need to address the “black box” nature of many complex models [2], as users often find it challenging to comprehend why an AI system produces a specific output, such as an item or product recommendation provided by a Recommender System. This lack of transparency has proved to lead to a diminished user experience, reduced trust from society in AI systems, and, in the worst case, potentially harmful outcomes [3]. This has popularly led to regulatory efforts to ensure transparency and accountability in AI systems. For instance, the European Union’s General Data Protection Regulation (GDPR) mandates the right to explanation in data-treating systems, granting users insight into any automated decisions made by them [4]. Ethical-oriented guidelines, such as those published by the EU’s High-Level Expert Group on Artificial Intelligence, emphasize transparency and accountability in AI [5]. Additionally, the U.S. Federal Trade Commission (FTC) emphasized the need for explainable AI to prevent deceptive practices in a recent report [6]. More recently, AI-specific legislation such as the EU’S AI Act has mandated AI systems to be designed in developed in such a way that their operator is sufficiently transparent to enable developers and users to interpret the system’s output and use it appropriately [7, 8].
@@
With respect to recommendations, the increasing reliance of businesses and users on AI-based Recommender Systems (RS) means that understanding and interpreting the recommendations provided to users has become crucial. Users expect personalized recommendations that align with their preferences, but they also desire explanations for why certain recommendations are made, e.g. due to a lack of quality of the recommendation or mistrust in the system; as such, explaining the outputs of RS not only enhances user trust, but also further enables users to make informed decisions, promoting a more interactive and engaging user experience [9].
@@
In the quest for explainability, one promising avenue is leveraging the visual content associated with AI inputs and outputs. Visual information, such as images, plays a vital role in human perception and decision-making processes; it has been explored before as a tool for explainability in other tasks [10]. In the context of user-to-item recommendations, there exist many examples of visual-based explainability: Netflix uses a small selection of movie thumbnails crafted by experts to cater to different types of users [11]; Chen, Xu, et al. [12] leverage user tastes from textual fashion reviews to highlight different image sections of the fashion item being recommended; Guesmi et al. [13] and Xu et al. [14] generate word cloud images to visually explain what user or item characteristics prompted article or movie recommendations. Consequently, utilizing visual content as a means to explain recommendations can greatly enhance user understanding and satisfaction [15] or the effectiveness of industrial processes [16].
@@
Within the idea of visual explanations, our particular approach of interest is the use of user-uploaded images to provide the explanations, avoiding a need for auxiliary information (e.g. text) to distil user tastes or a manual selection of creation of the images, thus yielding more organic, natural explanations and improving the self-sustainability and scalability of the explainable methodology. Compared to other visual-based approaches, which involve the generation of synthetic images, utilizing existing real images of the item recommended also provides a more implicit, self-contained and trustworthy explanation inside the recommendation pipeline, as the explanations will always be representative of the item’s features.
@@
Leveraging user-uploaded images is a promising but particularly challenging task: in the case of textual contents, a textual explanation may be associated with multiple users or items (e.g. the phrase “great visual effects” in movie reviews), which can help in tackling data sparsity [17]. User-uploaded visual explanations (e.g. an image taken in a restaurant), however, are unique to each user and item, making it difficult for a model to distil user tastes. This challenge has been tackled by ELVis [18] and MF-ELVis [19], which model and solve the explanation problem as a Learning to Rank (LTR) authorship task, predicting the best explanations of a (user, item) recommendation as a ranking of the images of the item most likely to have been authored by the user. Figure 1 illustrates this approach, which is also the object of research of this work.
Figure 1: Overview of the use of user-uploaded images for visual-based explainability and personalization recommendation. Users upload images to justify their experiences with an item, so image features are good explanations for the user; a model can learn these features for each user (left half). Then, the appeareance of any entering recommendation can be personalized (here, with an image of the recommended restaurant) using the item’s image that best reflects the learnt explanatory preferences of the user; the chosen image explanation will vary between users depending on their preferences (right side).
@@
While explainability is crucial, it should not come at the expense of sustainability and model compactness. AI and Deep Learning models are known for their high computational cost, carbon emissions and model sizes; the concept of sustainable or “green” AI has become a popular research topic in the past years [20]. These concerns have highlighted the need to balance model efficiency and explainability [21] [22]. ELVis first provided explainability to an RS by leveraging user-uploaded images, but used a Deep Learning-oriented architecture with a cost comparable to the RS itself. Although MF-Elvis optimized the architecture of the model with a simpler yet adequate Matrix Factorization approach, both models still need a costly training process and a large model size to obtain good results. We believe one potential culprit of this computational cost is the inadequacy of the training objective used by both ELVis and MF-ELVis, which model the task of ranking the best image explanations as a surrogate binary classification of image authorship, which may be far from optimal.
@@
    We introduce, design and implement BRIE, a novel model that leverages existing user-uploaded images to provide visually interpretable explanations for Recommender System outputs. BRIE improves existing approaches by adopting a more adequate learning objective closer to the final use of the model, while maintaining a highly optimized and simple architecture.
@@ 2.1 Problem Formulation
Therefore, leveraging user-uploaded item images as visual explanations poses two main challenges: 1) extracting meaningful latent information from photographs, given their limited association with multiple users and items, and 2) addressing data sparsity in real-world datasets where the available set of photographs $\mathcal{P}$ may be insufficient compared to the set of users $\mathcal{U}$, and no additional information (e.g. demographic variables) is available.
@@
As an explainability task, this problem assumes that the item $i$ being recommended is already given in a previous step by a Recommender System or any other information filtering system that matches users with items, and therefore focuses on finding a ranking function $f$ that can capture nuanced user preferences and photograph characteristics to provide adequate explanations of any given user-item pair $(u,i)$ while remaining disentangled from the overarching user-item pairing system.
@@ 2.2 ELVis and MF-ELVis
ELVis and MF-ELVis are the two existing models current state-of-the-art approaches for visual explainability through user-uploaded images, a novel approach for explainability in Recommender Systems. In this subsection, we describe the functioning of these models and provide insights into their design choices. Fig. 2 depicts in detail the network topologies of both models.
@@
Figure 2: Network topologies of ELVis and MF-ELVis, the two existing models for visual explainability through user-uploaded images in Recommender Systems. Both models take as inputs the IDs of user $u$ and photograph $p$, and output the predicted authorship $\mathbf{\hat{R}}_{up}$, i.e. how good is $p$ as an explanation of a recommendation of its associated item for user $u$. During training, both models minimize BCE loss $\mathcal{L}_{\text{BCE}}$ based on the predicted and real authorship $\mathbf{\hat{R}}_{up}$ and $\mathbf{R}_{up}$.
@@ 2.2.1 ELVis
Díez et al. [18] propose in ELVis to capture the underlying patterns and preferences in the interactions between users and photographs by representing the interaction matrix $\mathbf{R}$ as two latent matrices: the matrix of latent user preferences $\mathbf{U}$ and the matrix of latent photograph features $\mathbf{V^{\prime}}$.
ELVis generates a latent user vector $\mathbf{U}_{u}\in\mathbb{R}^{d}$ based on the user’s ID, and obtains a latent photograph vector $\mathbf{V}_{p}^{\prime}\in\mathbb{R}^{d}$ by projecting the ResNet-V2 [23] embedding $\mathbf{V}_{p}\in\mathbb{R}^{1536}$ of the original photograph into the latent factor subspace. The authors then propose to concatenate both vectors as $\mathbf{x}_{up}$, which is then passed through a multi-layer perceptron (MLP) $g$ to apply a series of non-linear transformations, such that
@@
The rationale behind ELVis’s design is based on using a learned similarity function, following the popular Neural Collaborative Filtering architecture [24]. The principle is that MLPs are general function approximators, so they should be -in theory- equally or more expressive than a fixed similarity function, such as a cosine similarity or dot product, and should capture more nuanced latent user preferences. However, the usage of an MLP comes at the expense of a very high computational cost overhead in both training and inference, motivating later approaches to abandon the usage of deep architectures.
@@ 2.2.2 MF-ELVis
A later publication suggests MF-ELVis [19] to simplify the model architecture while maintaining performance. Similar to ELVis, MF-ELVis learns the latent user preference matrix $\mathbf{U}$ and item feature matrix $\mathbf{V^{\prime}}$. However, instead of an MLP, MF-ELVis uses a fixed similarity function, the inner product, to compute the similarity between the corresponding user vector $\mathbf{U}_{u}$ and photograph vector $\mathbf{V}_{p}^{\prime}$. For a given (user, photograph) pair, the preference score is calculated as
@@
The rationale for adopting the inner product in MF-ELVis is based on findings presented by Rendle et al. [25]. The study suggests that, in similarity-related tasks, the classic inner product can achieve performance comparable to an MLP while simplifying the model architecture. Leveraging the inner product operation, MF-ELVis enhanced the time and carbon emissions required to train the model by simplifying the model architecture. However, it utilized a much higher number of latent factors $d$ to maintain competitiveness with ELVis, increasing the overall model size; this suggested that under the training protocols ELVis and MF-ELVis use, the model architecture could not be further simplified without losing performance, but other parts of the training pipeline, such as the modelled learning objective and the use of training data, could be suboptimal and an aspect to be improved.
@@ 2.3 Limitations of ELVis and MF-ELVis
Both ELVis and MF-ELVis address this issue by employing a uniform negative sampling and a positive sample replication, as shown in the left half of Figure 3, yielding a static training dataset 40 times larger than the original with a balanced number of positive and negative samples. Specifically, for each positive sample $(u,i,p)\in\mathcal{D}$, 39 additional samples are generated and used during all the training: 19 extra replications of $(u,i,p)$, 10 negative triads $(u,i,p^{\prime})$ (where $p^{\prime}$ are from the same restaurant as $p$ but uploaded by a different user $u^{\prime}$), and 10 negative triads $(u,i^{\prime},p^{\prime})$ (where $p^{\prime}$ were uploaded by a different user $u^{\prime}$ in a restaurant other than $i$). In the field of Positive-Unlabelled Learning, this random selection of negative samples is considered a naive yet straightforward method [26].
@@ 3 BRIE
In this section, we present BRIE (Bayesian Ranking of Images for Explainability), a novel model that builds upon MF-ELVis and addresses its limitations to surpass state-of-the-art performance in serving user-uploaded visual content to explain recommendations, while drastically improving training and inference efficiency.
@@ 3.1 Modelling Explanation Ranking with BPR
State-of-the-art approaches, like ELVis and MF-ELVis, utilize binary classification as a surrogate task to model the Learning to Rank scenario required to select the most adequate visual explanation of an item for a given user. Instead, BRIE uses the Bayesian Pairwise Ranking (BPR) [27] algorithm to learn the ranking function $f$. BPR is tailored to optimize model weights by maximizing the probability of correctly ranking positive “good” instances higher than negative “bad” instances.
The assumption for defining what is a positive example (“good instances”) and a negative example (“bad instances”) is similar to the one defined by ELVis and MF-ELVis:
  • 1.
    Regarding the positive examples, because the images of items uploaded by a given user $u$ (i.e. $\mathcal{P}_{u}$) are representative of what $u$ uses to justify and personalize its reviews, then all images $p\in\mathcal{P}_{u}$ can be considered positive examples for that user. For any user $u$, we also can reasonably assume smoothness in the characteristics of images inside $\mathcal{P}_{u}$: across different restaurants and reviews, the user’s images will generally focus on the same feature(s) of the items (e.g. a user that likes to upload images the interior design in each restaurant to justify their reviews).
  • 2.
    Concerning the negative examples, we maintain the simple, naive assumption made by ELVis and MF-ELVis, to better focalize the contributions of this work: any image not originally uploaded by $u$ may be selectable as a negative example for that user. For the sake of simplicity, this deliberately ignores that other images not originally uploaded by the user may share the explanatory features of the user’s own images, i.e. they could be possible positive examples for $u$. As we introduced in Section 2.3, this situation is common to data types where no real negative examples are known, so the full design of a PU Learning [26] technique to tackle this complex issue is left as future work.
Although we hold similar assumptions for the labelling training examples, the policy for sampling training examples in BRIE differs from the one ELVis and MF-ELVis utilize for their binary classification learning, described in Section 2.3. At the start of each training epoch, BRIE constructs an “extended” training set $\mathcal{D^{+}}={(u,i,p,p_{neg})}$ by uniformly sampling, for each $(u,i,p)\in\mathcal{D}$, a photograph $p_{neg}$ that is not part of $\mathcal{P}_{u}\subset\mathcal{P}$, i.e. a photograph from a purely random item and not taken by user $u$.
Our method is simpler than the pre-training selection used by ELVis and MF-ELVis, allows the regeneration of negative samples between epochs, and does not require storing all the pre-selected negative samples and replicated positive samples, unlike ELVis and MF-ELVis. Figure 3 compares this selection policy with the one used in current state-of-the-art models.
@@ 4.1 Datasets
The datasets used in our experiments are derived from the original publication of ELVis by Diez et al. [18], comprising photographs uploaded by users in the context of restaurant reviews on TripAdvisor, a popular hospitality establishment reviewing website. As far as our knowledge goes, this is the most diverse suite of datasets available for evaluating the use of user-uploaded images for visual-based RS explainability. We utilize public, real-world datasets from six different cities: Gijón, Barcelona, Madrid, New York, Paris, and London, in increasing number of photographs in the dataset, providing diverse contexts for evaluation in terms of socio-cultural characteristics, scale and user behaviours.
Each sample in the raw dataset consists of a triplet $(u,i,p)\in\mathcal{D}$, where $u$ represents the user ID, $i$ represents the restaurant ID, and $p$ represents the photograph ID. Table 1 summarizes key characteristics of each dataset, including the number of users, photographs, restaurants, and the overall sparsity of the datasets.
@@
For the sake of result reproducibility and comparability, we use the original train/validation/test partitions provided by the original authors. For each test case, we have one positive sample: a photograph $p$ that was not present in the training data but actually belongs to the user $u$. The remaining photographs $p^{\prime}\in(\mathcal{P}_{i}\setminus\mathcal{P}_{u}$) of the same restaurant, contributed by different users, serve as negative samples. The goal of the model is to rank the positive sample (the user’s own photograph) as high as possible among the negative samples, indicating its relevance as an explanation; this is a classic, popular approach to perform off-line evaluation in tasks involving implicit user-item interaction data [28]. Table 2 shows an example partitioning of each of the datasets used in this work.
Table 2: Example statistics (number of unique users and restaurants, and number of photographs) of a train-validation-test partitioning for each of the used datasets. Here, the number of photographs represents the number of positive examples in the partition, as well as the number of validation/test cases in the case of the validation/test sets.
Train Validation Test
Users Restaurants Photographs Users Restaurants Photographs Users Restaurants Photographs
Gijón 5,139 598 15,190 428 205 1,112 1,023 346 2,337
Barcelona 33,537 5,881 120,221 4,363 2,033 10,453 8,697 3,211 19,742
Madrid 43,268 6,810 162,487 5,852 2,301 14,276 11,874 3,643 27,142
New York 61,019 7,588 179,093 7,687 2,571 17,222 16,842 4,135 34,826
Paris 61,391 11,982 203,122 7,417 3,969 16,466 15,242 6,345 32,048
London 134,816 13,888 384,528 14,134 5,121 31,828 30,393 8,097 63,442
It is worth noticing that, as datasets do not have any explicit information about each user (e.g. demographic), any model must learn the user’s latent features before being able to select personalized explanations to that user (the classic “cold start” situation of user-item interaction data). To ensure a correct evaluation on this setting, Díez et al. [18] utilize a partitioning method common in tasks that deal with user-item interaction data [29], which ensures the positive example hidden in each test case isn’t a from new (cold-start) user, i.e. that all users are present in the train set; we refer readers to ELVIs’ original publication [18] for more detailed information on the dataset extraction, characteristics, and partitioning.